day 8 Aug 99

Marj Corcoran (corcoran@physics.rice.edu)
Sun, 8 Aug 1999 15:34:00 -0500 (CDT)

Shift Summary for DAY shift, 8 Aug 99

On shift: Theo, Marj
Also present: Julie, Ron, Rick K

Hours of beam available at all intensities: 7.5/8.5
Hours of beam available at nominal intensity:7.5/8.5
Hours of E832 data (special runs not included): 2.25
Hours of special runs (laser scans, muon runs...):none

Accounting of lost time: Problems with several CsI channels; problem
with keytree
==================================================================

Runs ended this shift:
14133-14149 Cosmic runs for CsI debugging
14150--good E832 data, in progress

========================================================================
Accesses during shift:
Many accesses for CsI problems.

========================================================================
Overview of shift:

Started shift with interlocks dropped and Ron, Rick, and Julie
working on CsI. Channels 1137 and 1542 are fixed, but channel 1064
shows up bad in laser illumination. Ron and Julie go back down and
swap divider. This channel is OK, but then 1136 dynode is dead. Access to
reseat it is successful, but bad capid's show up and some gains have
drifted. Notice rib temps are high; valve to balloons is not closed.
Ron and Julie close it, temps drop, bad capid's go away, gains stablize.
We start a run but channel 1539 appears to have a dead dynode,
one more access to reseat it. We are finally ready for beam, but the
keytree won't make up and interlock experts have to come in. We finally start
taking data at 13:15.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
14143 7:30 Cosmic run for debugging CsI channels 1542 and 1137.
Channels that were swapped were OK, but
channel 1064 shows up dead! Ron and Julie
head back downstairs, swap divider.
14144 8:20 Cosmic run to check CsI. All channels are live,
but pmt gain has drifted in laser plot

8:30 MCR called to do search and secure

14145 8:35 Cosmic run to look at channels which
have drifted.

9:15 Search and secure done, but Julie and Ron
notice a dynode (1136)is dead, controlled
access to reseat it, fix is successful.

9:45 Three channels show bad capid

10:10 Notice rib temps are rising. Ron and Julie go
back down and notice pressure relief valve for
balloons was not completely closed. Temps start
to drop as soon as it is closed.

14148 10:45 Rib temps back to normal. Start new cosmic run to
get laser illumination gains, reset defaults on
laser illum. gain plot.

14149 11:00 Start Beam_832 run w/o beam to check pmt gain
plot from quiescent laser. PMT gain plot looks
great, ask for beam.

14149 1-8 11:15 Beam stop was in for first 8 spills
14149 9 11:20 Beam stop out, but channel 1539 immediately
shows up as inefficient in the HCC map.

14149 20 11:40 Julie and Theo back in to look at ch 1539, reseat

11:50 Ch 1539 dynode looks OK on scope, end run,
14150 1 11:55 start new run

14150 16 12:15 Key tree won't make up properly, shutter won't
come up. MCR operators on the way. Shift crew
labels 40 tapes.

14150 61 13:15 Interlock guys fix keytree with a couple of
4-40 nuts! Beam returns, we are taking data!

14150 13:55 Jason H has checked the e/p for channels which
were swapped...all values w/in 1% of 1.0, so
they are OK

===================================================================

Followup notes on previous shift:

===================================================================
Items to follow up on next shift:

Keep an eye on the e/p for channels 1064,1137,1542.